OpenStudy (anonymous):

What is the amplitude, period, and phase shift of f(x) = −3 cos(4x + π) + 6?

OpenStudy (anonymous):

Y = Asin(Bx-C)+D A = amplitude 2pi/B = period C/B = phase shift

OpenStudy (anonymous):

I've tried solving and i keep getting confused

OpenStudy (anonymous):

amplitude = 3; period = pi over two; phase shift: x = negative pi over four amplitude = −3; period = pi over two; phase shift: x = pi over four amplitude = −3; period = negative pi over two; phase shift: x = negative pi over four amplitude = 3; period = 2π; phase shift: x = negative pi over four these are my options

OpenStudy (anonymous):

what did you get when you solved it?

OpenStudy (anonymous):

i think its either a or d?

OpenStudy (anonymous):

I'm thinking a

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Remember the amp is A the full A including the sign

OpenStudy (anonymous):

so I'm wrong?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so b ?

OpenStudy (anonymous):

lets work it one by one A = |-3| = 3 period = 2pi/4 = pi/2 phase shift = -pi/4

OpenStudy (anonymous):

it looks like its a

OpenStudy (anonymous):

so i was right.. not wrong. thnx

OpenStudy (anonymous):

yes